16/36 = [ ]/63 = 36/[ ] = [ ]/117
 16/36 = [ ]/63 = 36/[ ] = [ ]/117

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Consider the first two ratios 16/36 = [ ]/63 Let us assume the missing number be P Then, 16/36 = P/63

By cross multiplication we get,
P = (16 × 63)/36
P = 1008/36
Divide both numerator and denominator by 36.
P = 28

Therefore, 16/36 = [28]/63
Now, consider the middle two ratios, 28/63 = 36/[ ] Let us assume the missing number be Q,
Then, 28/63 = 36/Q
By cross multiplication we get,
Q = (36 × 63)/28
Q = 2268/28
Divide both numerator and denominator by 28. Q = 81
Therefore, 28/63 = 36/[81]
Consider the last two ratios 36/81 = [ ]/117
Let us assume the missing number be R
Then, 36/81 = R/117
By cross multiplication we get,
P = (36 × 117)/81
P = 4212/81
Divide both numerator and denominator by 81. P = 52
Therefore, 36/81 = [52]/117
So, 16/36 = [28]/63 = 36/[81] = [52]/117

Thanks